OpenStudy (anonymous):

Graph f (x) = x2 − 4 3x + 9

OpenStudy (anonymous):

Is it \[f(x)=(x^2-4)(3x+9)\]

OpenStudy (anonymous):

it is subtract

OpenStudy (anonymous):

not multiply

OpenStudy (cruffo):

\[f(x) = x^2 - 43x+9\]???

OpenStudy (anonymous):

f(x) = (x2-4)/(3x+9)

OpenStudy (anonymous):

It's very helpful if you use the equation editor to say specifically what you mean ;)

OpenStudy (anonymous):

f(x) = \[{x^2-4\over 3x+9}\]

OpenStudy (anonymous):

That it?

OpenStudy (anonymous):

yes

OpenStudy (cruffo):

Do you want to do this by hand or with a calculator?

OpenStudy (anonymous):

with by hand

OpenStudy (cruffo):

Have you found the vertical asymptote?

OpenStudy (anonymous):

I forgot a resolve

OpenStudy (cruffo):

I'm sorry, I don't understand.

OpenStudy (anonymous):

you can help me write the answer

OpenStudy (cruffo):

Ok, to graph this by hand you'll need to find the following: 1.) vertical asymptote(s) (if any) 2.) holes (if any) 3.) horizontal asymptote (if any) 4.) slant asymptote (if any) 5.) x- and y-intercepts 6.) some other points to help fill in the detail

OpenStudy (cruffo):

1.) to find the vertical asymptote you need to set the denominator equal to zero and solve for x. \[3x+9 = 0\]

OpenStudy (cruffo):

Post your answer when you get it so I can see you're on the right track :)

OpenStudy (anonymous):

x=-3

OpenStudy (cruffo):

great. We need to make sure that x=-3 is actually a VA and not a hole. To do this we need to substitute -3 into the numerator to see if we get 0. If we get 0 then x=-3 is a hole, if we don't get 0 then x=-3 is a VA. \[(-3)^2-4 =? \; 0\]

OpenStudy (cruffo):

So is it a hole?

OpenStudy (anonymous):

I can't find the answer

OpenStudy (cruffo):

What number do you get as an answer to the following: \[(-3)^2-4\]